The son is 21 years younger than the mother. In six years she will be five times as old as him. What is the father doing?
Found 2 solutions by richwmiller, mfisher333*:
Answer by richwmiller(17219)

m=s+21,
m+6=5*(s+6)
The father is doing 6 to 10 for multiple burglaries.

Answer by mfisher333*(1)

The father is impregnating the mother. 9 months is sons age mother is 20.25 in age.